OpenStudy (anonymous):

What is the second step to solve the equation twelve minus a equals seven ? Add 7 Add twelve Divide by negative one Subtract twelve

OpenStudy (anonymous):

\[12-a=7\]?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

you got to subtract to find -a

OpenStudy (anonymous):

there are many ways to solve this, so how can they ask "what is the second step"? question makes no sense

OpenStudy (anonymous):

one way would be \[12-a=7\] \[-a=7-12=-5\] \[a=5\] so i guess you could call the second step "divide by negative one" but that is rather silly

OpenStudy (anonymous):

and that is just for one method tell your teacher there is more than one way to skin a cat also "divide by negative one" is the same as "multiply by negative one" is the same as "change the sign of both sides" who wrote his???